During its most recent fiscal year, Raphael Enterprises sold 340,000 electric screwdrivers at a price of $19.20 each. Fixed costs amounted to $1,156,000 and pretax income was $1,496,000. What amount should have been reported as variable costs in the company's contribution margin income statement for the year in question?Multiple Choice

a. $5,032,000.

b. $2,652,000.

c. $3,876,000.

d. $6,528,000

.e. $2,720,000.
Answer #1

Total sales=(340,000*19.2)=$6,528,000

Total contribution margin=Fixed cost+Target profits

(1,156,000+1,496,000)=$2,652,000

Contribution margin=Sales-Variable cost  

Hence variable costs=6,528,000-2,652,000

=$3876000